The perimeter of a rectangle is 60 inches and the area is 161 square inches. Find the dimensions of the rectangle

Answer by rothauserc(4718)

area(A) = l * w where l is length and w is width
we are given l * w = 161, therefore
*****************************************************
we ask "what are the prime factors of 161"?
161's prime factors are 7 and 23, that is 7 * 23 = 161
******************************************************
Perimeter = 2l + 2w
let l = 23 and w = 7, then
60 = 46 + 14
60 = 60
therefore
the dimensions of the rectangle are
******************************************************
length = 23 inches and width = 7 inches